If the current in the 4.0- ohm resistor is 1.4 A, what is the magnitude of the potential difference, VA - VB?
1
Expert's answer
2016-04-23T08:42:06-0400
dV = I * R;

VA - VB = 1.4 * 4.0 = 5.6 V